Grade 128 grade maths

wat is 8x25+50-25/5 =?

mamam
11 Years agoGrade 12

8 Answers

Prajwal Kavad
ApprovedApproved Tutor Answer11 Years ago
245 is the answer
mamam
11 Years ago
thank you
V. Harshit -
11 Years ago
245 is the answer
 
Adarsh
11 Years ago
 8x25+50-25/5 =200+50-5=245 
Abishek arun
11 Years ago
According to the BODMAS rule
8x25+50-25/5 
first divide
8x25+50-25/5
8x25+50-5
secondly multiply
8x25+50-5
200+50-5 = 245
BODMAS rule follows as
brackets then off then divide  then multiply then add then subtract
hope you understand 
please approve my answer
deapan
11 Years ago
answer is 245 by following the BODMAS rule
please approve my answer
Hinal Srivastava
11 Years ago
 
According to the BODMAS rule
8x25+50-25/5 
first divide
8x25+50-25/5
8x25+50-5
secondly multiply
8x25+50-5
200+50-5 = 245
Hope it helps!!!
SEELA YUVARAJU
10 Years ago
245